The weight of newborn baby boys at a local hospital are believed to have a normal distribution with mean weight of 3943 grams and a variance of 51984. If a newborn baby boy born at hospital is randomly selected, find the probability that the weight will be between 4262 and 4353 grams. Round to 4
Answer #1

Solution :

Given that ,

mean = \mu = 3943

Variance = 51984

standard deviation = \sigma = 228

P(4262 < x < 4353) = P[(4262 - 3943)/228 ) < (x - \mu ) /\sigma  < (4353 - 3943) /228 ) ]

= P(1.40 < z < 1.80)

= P(z < 1.80) - P(z < 1.40)

= 0.9641 - 0.9192

0.0449

Probability = 0.0449